# Assessing Omitted Variable Bias

### v5

@@ Abstract
Omitted variables are one of the most important threats to the identification of causal effects. Several widely used methods assess the impact of omitted variables on empirical conclusions by comparing measures of selection on observables with measures of selection on unobservables. The recent literature has discussed various limitations of these existing methods, however. This includes a companion paper of ours which explains issues that arise when the omitted variables are endogenous, meaning that they are correlated with the included controls. In the present paper, we develop a new approach to sensitivity analysis that avoids those limitations, while still allowing researchers to calibrate sensitivity parameters by comparing the magnitude of selection on observables with the magnitude of selection on unobservables as in previous methods. We illustrate our results in an empirical study of the effect of historical American frontier life on modern cultural beliefs. Finally, we implement these methods in the companion Stata module regsensitivity for easy use in practice.
@@
Keywords: Treatment Effects, Sensitivity Analysis, Unconfoundedness
@@ 1 Introduction
For many years, researchers’ best approach to dealing with the threat of omitted variables was to either sign the bias, or to use a different identification strategy altogether. But oftentimes no alternative identification strategies are available, and the direction of the bias works against them—meaning that plausible omitted variables could explain away their results. In response to this challenge, the literature developed a variety of results to quantify the impact of omitted variables. A particularly important breakthrough was the paper by Altonji et al. (2005), and later extended by Oster (2019), who showed how to measure the impact of omitted variables by comparing the magnitude of selection on observed variables with the magnitude of selection on unobserved variables. The tools developed in these papers are now widely used in empirical economics.
However, the recent statistics and econometrics literature has discussed several limitations with those early, first generation methods. This includes De Luca et al. (2019), Cinelli and Hazlett (2020, section 6.3), Basu (2022), Masten and Poirier (2023), and a companion paper of ours, Diegert et al. (2025). In particular, in our companion paper Diegert et al. (2025), we show that previous methods can lead to incorrect conclusions about robustness when the controls are endogenous, meaning that they are correlated with the omitted variables of concern. For example, consider a classic regression of wages on education and controls like parental education. Typically we are worried that the coefficient on education in this regression is a biased measure of the returns to schooling because unobserved ability is omitted. In this example, the controls would be exogenous if unobserved ability is uncorrelated with parents’ education, along with all other included controls. Such exogeneity assumptions on the control variables are usually thought to be very strong and implausible. For example, Angrist and Pischke (2017) discuss how
@@
Consequently, it is important that methods for assessing sensitivity to omitted variables allow the controls to be endogenous. In this paper, we develop one such method, based on a new set of sensitivity parameters. Like the previous methods, these new parameters still allow researchers to measure the importance of unobserved variables by comparing them to the included covariates. Thus researchers can still measure how strong selection on unobservables must be relative to selection on observables in order to overturn their baseline findings, even if the controls are endogenous. Moreover, as proved in our companion paper Diegert et al. (2025), this sensitivity parameter does not suffer from one of the main critiques of previous methods. This allows it to be used even in applications with endogenous controls.
@@
In section 2 we describe the baseline model. The parameter of interest is $\beta_{\text{long}}$, the coefficient on a treatment variable $X$ in a long OLS regression of an outcome variable $Y$ on a constant, treatment $X$, the observed covariates $W_{1}$, and some unobserved covariates $W_{2}$. In section 4 we discuss causal models which allow us to interpret this parameter causally, based on three different identification strategies: unconfoundedness, difference-in-differences, and instrumental variables. Since $W_{2}$ is unobserved, we cannot compute the long regression of $Y$ on $(1,X,W_{1},W_{2})$ in the data. Instead, we can only compute the medium regression of $Y$ on $(1,X,W_{1})$. We begin by considering a baseline model with a “no selection on unobservables” assumption, which implies that the coefficients on $X$ in the long and medium regressions are the same. Importantly, this baseline model does not assume the controls $W_{1}$ are exogenous. They are included solely to aid in the identification of the coefficient on $X$ in the long regression.
In section 3 we develop new methods for assessing sensitivity to omitted variables that (a) do not rely on the exogenous controls assumption, (b) satisfy the consistency and monotonicity-in-selection criteria that we develop in our companion paper Diegert et al. (2025), and (c) still allow researchers to compare the magnitude of selection on observables with the magnitude of selection on unobservables. Our identification results use either one or two sensitivity parameters. The first sensitivity parameter compares the relative magnitude of the coefficients on the observed and unobserved covariates in a treatment selection equation. This parameter thus measures the magnitude of selection on unobservables by comparing it with the magnitude of selection on observables. The second sensitivity parameter compares the relative magnitude of the coefficients on the observed and unobserved covariates in the outcome equation. In section 3.5, we extend our results to allow researchers to also restrict the magnitude of control endogeneity.
@@
If researchers are willing to restrict the impact of unobservables on outcomes, then they can obtain results that are more robust to selection on unobservables. In this case, the identified set is more difficult to characterize analytically (see Theorem LABEL:thm:mainMultivariateResult in the appendix). However, our second main identification result (Theorem 3) shows that we can nonetheless easily numerically compute breakdown points. We generalize this result in section 3.5 to also allow restrictions on the endogeneity of the controls.
In section 5 we show how to use our results in empirical practice. We use data from Bazzi et al. (2020, Econometrica) who studied the effect of historical American frontier life on modern cultural beliefs. Specifically, they test a well known conjecture that living on the American frontier cultivated individualism and antipathy to government intervention. Among other analyses, they use Oster’s (2019) method to argue against the importance of omitted variables. Using our results, we obtain richer conclusions about robustness. In particular, when allowing for endogenous controls, we find that effects obtained from questionnaire based outcomes are often sensitive to omitted variables while the effects from election and property tax outcomes remain robust.
@@
Although control variables are often thought to be endogenous, the literature on sensitivity analysis generally either assumes the controls are exogenous or uses a method called “residualization.” This includes Altonji, Elder, and Taber (2005, 2008), Krauth (2016), and Oster (2019). Imbens (2003) starts from the standard unconfoundedness assumption which allows endogenous controls, but in his parametric assumptions (see his likelihood equation on page 128) he assumes that the unobserved omitted variable is independent of the observed covariates. Altonji et al. (2019) propose an approach to allow for endogenous controls based on imposing a factor model on all covariates, observable and unobservable. Their approach and ours impose different assumptions and therefore are not nested. Cinelli and Hazlett (2020) develop a regression sensitivity analysis based on judgments about the R-squared from the long regression of treatment on the observed and unobserved variables. Their results do not rely on endogenous controls, but also do not allow researchers to compare selection on observables with unobservables in the tradition of Altonji et al. (2005) and Oster (2019).
@@
Notation Remark
For random vectors $A$ and $B$, let $\operatorname*{cov}(A,B)$ be the $\text{dim}(A)\times\text{dim}(B)$ matrix whose $(i,j)$th element is $\operatorname*{cov}(A_{i},B_{j})$. Define $A^{\perp B}\coloneqq A-\operatorname*{cov}(A,B)\operatorname*{var}(B)^{-1}B$. For a random variable $A$, let $R_{A\sim B\mathrel{{\bullet}}C}^{2}$ denote the R-squared from a regression of $A^{\perp C}$ on $(1,B^{\perp C})$.
@@ 2 The Baseline Model
The setup here is similar to our companion paper Diegert et al. (2025). Let $Y$ and $X$ be observed scalar variables. Let $W_{1}$ be a vector of observed covariates of dimension $d_{1}$ and $W_{2}$ be an unobserved vector of dimension $d_{2}$. Let $W=(W_{1},W_{2})$. Consider the population linear regression of $Y$ on $X$, $W_{1}$, $W_{2}$, and a constant. Let $\beta_{\text{long}}$ denote the coefficient on $X$, while $(\gamma_{1},\gamma_{2})$ denote the coefficients on $(W_{1},W_{2})$. We make the following assumption to ensure these linear regression coefficients are well-defined.
@@
This theorem shows why assumption A2 can also be interpreted as a “selection on observables” assumption: adjusting linearly for the observable $W_{1}$ is sufficient to recover $\beta_{\text{long}}$. Note that this result allows for endogenous controls, in the sense that the observed covariates $W_{1}$ can be arbitrarily correlated with the unobserved covariates $W_{2}$. But it restricts the relationship between $(X,W_{1},W_{2})$ in such a way that we can still point identify $\beta_{\text{long}}$ even though $W_{1}$ and $W_{2}$ are arbitrarily correlated. The coefficient $\gamma_{1}$ on the observed controls, however, is completely unidentified.
In practice, we are often worried that the no selection on unobservables assumption A2 does not hold. For example, consider the regression of wages on education discussed in section 1. Even with a rich set of observed covariates $W_{1}$, like family background variables, researchers typically suspect that there may still be omitted variables that affect both wages and education, and which could be correlated with the observed variables like family background. We therefore develop a new approach to assess the importance of this no selection on unobservables assumption in the next section.
@@ 3 Sensitivity Analysis
There are many ways to relax assumption A2 ($\pi_{2}=0$) and allow for some selection on unobservables. A key challenge is to construct a measure of selection on unobservables which is interpretable, and therefore allows researchers to reason about the magnitudes of this measure. In this section we propose a new measure of selection, argue for its interpretability, and then show how to use it to perform a sensitivity analysis to the selection on observables assumption.
3.1 The Relative Variation in Selection Equation Indices
Recall the regression of treatment on a constant and all the covariates:
@@ 3.1 The Relative Variation in Selection Equation Indices
Both the numerator and denominator are invariant to invertible linear transformations of $W_{1}$ or $W_{2}$, including rescalings, since the indices $\pi_{1}^{\prime}W_{1}$ and $\pi_{2}^{\prime}W_{2}$ are both invariant with respect to invertible linear transformations. This invariance ensures that $r_{X}$ is a unit-free measure of the relative magnitude of selection on unobservables to selection on observables. $r_{X}=0$ corresponds to baseline assumption A2 ($\pi_{2}=0$) while $r_{X}>0$ allows for some selection on unobservables.
@@
3.2 Identification Using Treatment Equation Assumptions Only
Recall from section 2 that our parameter of interest is $\beta_{\text{long}}$, the OLS coefficient on $X$ in the long regression of $Y$ on $(1,X,W_{1},W_{2})$. Since $W_{2}$ is not observed, we cannot compute this regression from the data. Instead, we can compute $\beta_{\text{med}}$, the OLS coefficient on $X$ in the medium regression of $(1,X,W_{1})$. The difference between these two regression coefficients is given by the classic omitted variable bias formula. To bound this difference, we restrict the selection ratio $r_{X}$ described in section 3.1. Specifically, consider the following assumption.
@@ 3.2 Identification Using Treatment Equation Assumptions Only
Let $\mathcal{B}_{I}(\bar{r}_{X})$ denote the identified set for $\beta_{\text{long}}$ under the positive definite variance assumption A1 and the restriction A3 on $\pi_{2}$. In particular, this identified set does not impose any restrictions on the impact of omitted variables on the outcome variable. Let
@@
$\operatorname*{cov}(W_{1},Y)\neq\operatorname*{cov}(W_{1},X)\operatorname*{cov}(X,Y)$ and $\operatorname*{cov}(W_{1},X)\neq 0$.
This assumption is not necessary, but it simplifies the proofs. A sufficient condition for A4 is $\beta_{\text{short}}\neq\beta_{\text{med}}$, where $\beta_{\text{short}}$ is the coefficient on $X$ in the short OLS regression of $Y$ on $(1,X)$. For simplicity, we also normalize the treatment variable so that $\operatorname*{var}(X)=1$ and the covariates so that $\operatorname*{var}(W_{1})=I$. All of the results below can be rewritten without these normalizations at the cost of additional notation.
@@
Suppose the joint distribution of $(Y,X,W_{1})$ is known. Suppose A1, A3, and A4 hold. Normalize $\operatorname*{var}(X)=1$ and $\operatorname*{var}(W_{1})=I$. If $\bar{r}_{X}<\sqrt{1-R^{2}_{X\sim W_{1}}}$, then
@@
Appendix LABEL:sec:mainResultProofIntuition gives a brief sketch derivation of this result, while the complete formal derivation and sharpness proof is in appendix LABEL:sec:proofsForMainNewAnalysis. Theorem 2 characterizes the largest and smallest possible values of $\beta_{\text{long}}$ when some selection on unobservables is allowed, the observed covariates are allowed to be arbitrarily correlated with the unobserved covariate, and we make no restrictions on the coefficients in the outcome equation. In fact, we prove that, with the exception of at most three singletons, the interval $[\underline{B}(\bar{r}_{X}),\overline{B}(\bar{r}_{X})]$ is the identified set for $\beta_{\text{long}}$ under these assumptions. Here we focus on the smallest and largest elements to avoid technical digressions that are unimportant for applications.
There are two important features of Theorem 2: First, it only requires researchers to reason about one sensitivity parameter, unlike some existing approaches. Second, it allows for arbitrarily endogenous controls since it imposes no restrictions on the association between $W_{1}$ and $W_{2}$. So this result allows researchers to examine the impact of selection on unobservables on their baseline results without also having to reason about the magnitude of endogenous controls.
Since Theorem 2 provides explicit expressions for the bounds, we can immediately derive a few of their properties. First, when $\bar{r}_{X}=0$, the bounds collapse to $\beta_{\text{med}}$, the point estimand from the baseline model with no selection on unobservables. Thus, we recover the baseline model as a special case. For small values of $\bar{r}_{X}>0$, the bounds are no longer a singleton, but their width increases continuously as $\bar{r}_{X}$ increases away from zero. After normalizing treatment and the observed covariates, the rate at which the bounds increase depends on just a few features of the data: all can be derived from the variance matrix of $(Y,X,W_{1})$. We also see that the bounds are symmetric around $\beta_{\text{med}}$. Finally, the bounds are finite only if $\bar{r}_{X}<\sqrt{1-R^{2}_{X\sim W_{1}}}$.
@@
The breakdown point described in Corollary 1 characterizes the magnitude of selection on unobservables relative to selection on observables needed to overturn one’s baseline findings. Corollary 1 explicitly shows that this breakdown point depends on just two features of the observed data: The relationship between treatment and the outcome, after adjusting for the observed covariates, and the first stage relationship between treatment and the observed covariates. In particular, as the covariate adjusted relationship between outcomes and treatment strengthens, the breakdown point increases too. In contrast, as the relationship between treatment and covariates strengthens, the breakdown point decreases. This follows since we are using effects of the observed covariates to calibrate the magnitude of selection on unobservables. So when these observed covariates are strongly related to treatment, we need relatively less selection on unobservables to overturn our baseline findings. We discuss this point more in section 3.4.
Recommendations. First, we recommend that researchers present estimates of the breakdown point $\bar{r}_{X}^{\text{bp}}$ as a scalar measure of the robustness of their results. This can be displayed alongside the main result point estimates, as in table 5.2 in our empirical application. Second, we recommend that researchers present estimates of the bounds described in Theorem 2 as a function of $\bar{r}_{X}$. Researchers may also find it useful to present estimates of the calibration parameters $\rho_{k}$ that we describe in section 3.4. We illustrate all of these recommendations in our empirical application in section 5.

### v6

@@ Abstract
Omitted variables are one of the most important threats to the identification of causal effects. Several widely used methods assess the impact of omitted variables on empirical conclusions by comparing measures of selection on observables with measures of selection on unobservables. The recent literature has discussed various limitations of these existing methods, however. This includes challenges that arise when the omitted variables are endogenous, meaning that they are correlated with the included controls. We develop a new approach to regression sensitivity analysis that avoids those limitations, while still allowing researchers to calibrate sensitivity parameters by comparing the magnitude of selection on observables with the magnitude of selection on unobservables as in previous methods. We illustrate our results in an empirical study of the effect of historical American frontier life on modern cultural beliefs. Finally, we implement these methods in the companion Stata module regsensitivity for easy use in practice.
@@
Keywords: Linear Regression, Sensitivity Analysis, Treatment Effects, Unconfoundedness
@@ 1 Introduction
For many years, researchers’ best approach to dealing with the threat of omitted variables was to either sign the bias, or to use a different identification strategy altogether. These include, for example, measurement error models using proxies, instrumental variable models, or methods for combining datasets where some of the omitted variables are observed. But oftentimes no viable alternative identification strategies are available, and the direction of the bias works against them—meaning that plausible omitted variables could explain away their results. In response to this challenge, the literature developed a variety of results to quantify the impact of omitted variables. A particularly important breakthrough was the paper by Altonji et al. (2005), and later extended by Oster (2019), who showed how to measure the impact of omitted variables by comparing the magnitude of selection on observed variables with the magnitude of selection on unobserved variables. The tools developed in these papers are now widely used in empirical economics.
However, the recent statistics and econometrics literature has discussed several limitations with those early, first generation methods. This includes De Luca et al. (2019), Cinelli and Hazlett (2020, section 6.3), Basu (2022), Masten and Poirier (2025), and a companion paper of ours, Diegert et al. (2025). [^4] See section 5 of Masten and Poirier (2025) for a detailed review and explanation of all of these limitations. In particular, in our companion paper Diegert et al. (2025), we show that previous methods can lead to incorrect conclusions about robustness when the controls are endogenous, meaning that they are correlated with the omitted variables of concern. For example, consider a classic regression of wages on education and controls like parental education. Typically we are worried that the coefficient on education in this regression is a biased measure of the returns to schooling because unobserved ability is omitted. In this example, the controls would be exogenous if unobserved ability is uncorrelated with parents’ education, along with all other included controls. Such exogeneity assumptions on the control variables are usually thought to be very strong and implausible. For example, Angrist and Pischke (2017) discuss how
@@
In this paper, we develop a method for assessing sensitivity to omitted variables that allows the controls to be endogenous. Our approach uses a new set of sensitivity parameters that have several important features:
  1. 1.
    They are based a new and simple measure of selection applied to both the observables and unobservables. Similar to Altonji et al. (2005) and Oster (2019), we take ratios of these measures to compare selection on unobservables with selection on observables in a way that treats those variables symmetrically. Thus researchers can still use our results to measure how strong selection on unobservables must be relative to selection on observables in order to overturn their baseline findings, even if the controls are endogenous. This is the main contribution of our paper.
  2. 2.
    They have a natural causal interpretation as ratios of causal effects of covariates on either the treatment or the outcome.
  3. 3.
    They allow researchers to impose assumptions on the correlation between the observed controls and the omitted variables. Upper bounds on this correlation can be used to impose exogenous or partially exogenous controls. This allows us to assess the extent to which endogenous controls affect the robustness of one’s results. Lower bounds on this correlation can also be imposed to reflect situations where the observed controls are believed to be sufficiently endogenous.
  4. 4.
    They do not suffer from one of the main critiques of previous methods, as proved in our companion paper Diegert et al. (2025).
No other approaches in the literature have all of these features.
@@
In section 2 we describe the baseline model. The parameter of interest is $\beta_{\text{long}}$, the coefficient on a treatment variable $X$ in a long OLS regression of an outcome variable $Y$ on a constant, treatment $X$, the observed covariates $W_{1}$, and some unobserved covariates $W_{2}$. In section 4 we discuss causal models which allow us to interpret this parameter causally, based on three different identification strategies: unconfoundedness, difference-in-differences, and instrumental variables. Since $W_{2}$ is unobserved, we cannot compute the long regression of $Y$ on $(1,X,W_{1},W_{2})$ in the data. Instead, we can only compute the medium regression of $Y$ on $(1,X,W_{1})$. We begin by considering a baseline model with a “no selection on unobservables” assumption, which implies that the coefficients on $X$ in the long and medium regressions are the same. Importantly, this baseline model does not assume the controls $W_{1}$ are exogenous.
In section 3 we develop new methods for assessing sensitivity to omitted variables that (a) do not rely on the exogenous controls assumption, (b) satisfy the consistency and monotonicity-in-selection criteria that we develop in our companion paper Diegert et al. (2025), and (c) still allow researchers to compare the magnitude of selection on observables with the magnitude of selection on unobservables. Our identification results can be used with either one or two sensitivity parameters. The first sensitivity parameter compares the relative magnitude of the coefficients on the observed and unobserved covariates in a treatment selection equation. This parameter thus measures the magnitude of selection on unobservables by comparing it with the magnitude of selection on observables. The second sensitivity parameter compares the relative magnitude of the coefficients on the observed and unobserved covariates in the outcome equation. In section 3.5, we extend our results to allow researchers to also restrict the magnitude of control endogeneity.
@@
If researchers are willing to restrict the impact of unobservables on outcomes, then they can obtain results that are more robust to selection on unobservables. In this case, the identified set is more difficult to characterize analytically (see Theorem 5 in the appendix). However, our second main identification result (Theorem 2) shows that we can nonetheless easily numerically compute breakdown points. We generalize this result in section 3.5 to also allow restrictions on the endogeneity of the controls.
In section 5 we show how to use our results in empirical practice. We use data from Bazzi et al. (2020, Econometrica) who studied the effect of historical American frontier life on modern cultural beliefs. Specifically, they test a well known conjecture that living on the American frontier cultivated individualism and antipathy to government intervention. Among other analyses, they use Oster’s (2019) method to argue that omitted variables do not affect their conclusions. Using our results, we obtain richer conclusions about robustness. In particular, when allowing for endogenous controls, we find that effects obtained from questionnaire based outcomes are often sensitive to omitted variables while the effects from election and property tax outcomes remain robust.
@@
Although control variables are often thought to be endogenous, the literature on sensitivity analysis generally either assumes the controls are exogenous or uses a method called “residualization.” This includes Altonji, Elder, and Taber (2005, 2008), Krauth (2016), and Oster (2019). Imbens (2003) starts from the standard unconfoundedness assumption which allows endogenous controls, but in his parametric assumptions (see his likelihood equation on page 128) he assumes that the unobserved omitted variable is independent of the observed covariates. Altonji et al. (2019) propose an approach to allow for endogenous controls based on imposing a factor model on all covariates, observable and unobservable. Their approach and ours impose different assumptions and therefore are not nested. Cinelli and Hazlett (2020) develop a regression sensitivity analysis based on partial R-squareds. Their results do not rely on endogenous controls, but they do not define any measures of selection and do not work with symmetric ratio parameters like Altonji et al. (2005), Oster (2019), or us, and hence their results do not allow researchers to compare selection on observables with unobservables as in Altonji et al. (2005), Oster (2019), and our results.
@@
Notation Remarks
For random vectors $A$ and $B$, let $\operatorname*{cov}(A,B)$ be the $\text{dim}(A)\times\text{dim}(B)$ matrix whose $(i,j)$th element is $\operatorname*{cov}(A_{i},B_{j})$. Define $A^{\perp B}\coloneqq A-\operatorname*{cov}(A,B)\operatorname*{var}(B)^{-1}B$, the intercept plus the residual from a regression of $A$ on $(1,B)$. For a scalar random variable $A$, let $R_{A\sim B\mathrel{{\bullet}}C}^{2}$ denote the R-squared from a regression of $A^{\perp C}$ on $(1,B^{\perp C})$; this is called the partial R-squared (of $A$ on $B$, partialling out $C$).
@@ 2 The Baseline Model
Let $Y$ and $X$ be observed scalar variables. Let $W_{1}$ be a vector of observed covariates of dimension $d_{1}$ and $W_{2}$ be an unobserved vector of dimension $d_{2}$. Let $W\coloneqq(W_{1},W_{2})$. Consider the population linear regression of $Y$ on $X$, $W_{1}$, $W_{2}$, and a constant. Let $\beta_{\text{long}}$ denote the coefficient on $X$, while $(\gamma_{1},\gamma_{2})$ denote the coefficients on $(W_{1},W_{2})$. The setup here is similar to our companion paper Diegert et al. (2025). We make the following assumption to ensure these linear regression coefficients are well-defined.
@@
This result shows why Assumption A2 can also be interpreted as a “selection on observables” assumption: adjusting linearly for the observable $W_{1}$ is sufficient to recover $\beta_{\text{long}}$. Note that this result allows for endogenous controls, in the sense that the observed covariates $W_{1}$ can be arbitrarily correlated with the unobserved covariates $W_{2}$. But it restricts the relationship between $(X,W_{1},W_{2})$ in such a way that we can still point identify $\beta_{\text{long}}$ even though $W_{1}$ and $W_{2}$ are arbitrarily correlated. The coefficient $\gamma_{1}$ on the observed controls, however, is completely unidentified. Note that Proposition 1 is unsurprising and is not the main contribution of this paper; we include it for completeness because we have not seen it formally stated in any previous references, especially the second part that $\gamma_{1}$ is completely unidentified.
In practice, we are often worried that the no selection on unobservables assumption does not hold. For example, consider the regression of wages on education discussed in section 1. Even with a rich set of observed covariates $W_{1}$, like family background variables, researchers typically suspect that there may still be omitted variables that affect both wages and education, and which could be correlated with observed variables like family background. We therefore develop a new approach to assess the importance of this no selection on unobservables assumption in the next section.
@@ 3 Sensitivity Analysis
There are many ways to relax Assumption A2 ($\pi_{2}=0$) and allow for some selection on unobservables. A key challenge is to construct measures of selection on unobservables and observables that are interpretable and comparable, and therefore allow researchers to reason about their relative magnitudes. In this section we propose a new measure of selection, argue for its interpretability, and then show how to use it to perform a sensitivity analysis to the selection on observables assumption.
3.1 Measuring Selection
We begin by focusing on the relationship between treatment and the covariates, observed and unobserved. Later, in section 3.3, we will extend our analysis to also consider the relationship between outcomes and the covariates. Several previous approaches in the literature use a single sensitivity parameter to measure the combined impact of omitted variables on outcomes and treatment; this leads to a variety of issues and limitations, however, as discussed in the papers cited in section 1, and hence we instead consider these two relationships separately. To that end, recall the regression of treatment on a constant and all the covariates:
@@ 3.1 Measuring Selection
This is a relative measure that measures the impact of unobservables as a multiple of the impact of observables. Consequently, in models where the included covariates $W_{1}$ are very important for treatment selection even a small value of $r_{X}$ might be considered to be a large impact of the unobservables; we discuss this further in section 3.4.
Both the numerator and denominator of $r_{X}$ are invariant to invertible linear transformations of $W_{1}$ or $W_{2}$, including rescalings, since the indices $\pi_{1}^{\prime}W_{1}$ and $\pi_{2}^{\prime}W_{2}$ are both invariant with respect to invertible linear transformations. This invariance ensures that $r_{X}$ is a unit-free measure of the relative magnitude of selection on unobservables to selection on observables. $r_{X}=0$ corresponds to baseline assumption A2 ($\pi_{2}=0$) while $r_{X}>0$ allows for some selection on unobservables.
@@
3.2 Identification Using Selection Equation Assumptions Only
Recall from section 2 that our parameter of interest is $\beta_{\text{long}}$, the OLS coefficient on $X$ in the long regression of $Y$ on $(1,X,W_{1},W_{2})$. Since $W_{2}$ is not observed, we cannot compute this regression from the data. However, we can compute $\beta_{\text{med}}$, the OLS coefficient on $X$ in the medium regression of $(1,X,W_{1})$. The difference between these two regression coefficients is given by the classic omitted variable bias formula. To bound this difference, we restrict the selection ratio $r_{X}$ described in section 3.1. Specifically, consider the following assumption.
@@ 3.2 Identification Using Selection Equation Assumptions Only
Let $\mathcal{B}_{I}(\bar{r}_{X})$ denote the identified set for $\beta_{\text{long}}$ under the positive definite variance assumption A1 and the restriction A3. In particular, this identified set does not impose any restrictions on the impact of omitted variables on the outcome variable. Let
@@
$\operatorname*{cov}(W_{1},Y)\neq\operatorname*{cov}(W_{1},X)\operatorname*{cov}(X,Y)/\operatorname*{var}(X)$ and $\operatorname*{cov}(W_{1},X)\neq 0$.
This assumption is not necessary, but it simplifies the proofs. Moreover, this assumption usually holds in practice since it holds whenever $\beta_{\text{short}}\neq\beta_{\text{med}}$, where $\beta_{\text{short}}$ is the coefficient on $X$ in the short OLS regression of $Y$ on $(1,X)$.
@@
Suppose the joint distribution of $(Y,X,W_{1})$ is known. Suppose A1, A3, and A4 hold. If $\bar{r}_{X}<\sqrt{1-R^{2}_{X\sim W_{1}}}$, then
@@
Appendix A gives a sketch derivation of and intuition for this result, while the complete formal derivation and sharpness proof is in Appendix LABEL:sec:proofsForMainNewAnalysis. Theorem 1 characterizes the largest and smallest possible values of $\beta_{\text{long}}$ when some selection on unobservables is allowed, the observed covariates are allowed to be arbitrarily correlated with the unobserved covariate, and we make no restrictions on the coefficients in the outcome equation. In fact, we prove that, with the exception of at most three singletons, the interval $\underline{B}(\bar{r}_{X}),\overline{B}(\bar{r}_{X})]$ is the identified set for $\beta_{\text{long}}$ under these assumptions. Here we focus on the smallest and largest elements to avoid technical digressions that are unimportant for applications.
There are two important features of Theorem 1: First, it only requires researchers to reason about one sensitivity parameter, unlike some existing approaches (our generalizations in sections 3.3 and 3.5 use additional sensitivity parameters to allow researchers to impose further restrictions). Second, it allows for arbitrarily endogenous controls since it imposes no restrictions on the association between $W_{1}$ and $W_{2}$. So this result allows researchers to examine the impact of selection on unobservables on their baseline results without also having to reason about the magnitude of endogenous controls.
Since Theorem 1 provides explicit expressions for the bounds, we can immediately derive a few of their properties. First, when $\bar{r}_{X}=0$, the bounds collapse to $\beta_{\text{med}}$, the point estimand from the baseline model with no selection on unobservables. Thus, we recover the baseline model as a special case. For small values of $\bar{r}_{X}>0$, the bounds are no longer a singleton, but their width increases continuously as $\bar{r}_{X}$ increases away from zero. The rate at which the bounds increase depends on just a few features of the data: All can be derived from the variance matrix of $(Y,X,W_{1})$. We also see that the bounds are symmetric around $\beta_{\text{med}}$. Finally, the bounds are finite only if $\bar{r}_{X}<\sqrt{1-R^{2}_{X\sim W_{1}}}$.
@@
The breakdown point described in Corollary 1 characterizes the magnitude of selection on unobservables relative to selection on observables needed to overturn one’s baseline findings. Corollary 1 explicitly shows that this breakdown point depends on just two features of the observed data: The relationship between treatment and the outcome, after adjusting for the observed covariates, and the first stage relationship between treatment and the observed covariates. In particular, as the covariate adjusted relationship between outcomes and treatment strengthens, the breakdown point increases too. In contrast, as the relationship between treatment and covariates strengthens, the breakdown point decreases. This follows since we are using effects of the observed covariates to calibrate the magnitude of selection on unobservables. So when these observed covariates are strongly related to treatment, we need relatively less selection on unobservables to overturn our baseline findings. We further discuss this point in section 3.4.
Practical Recommendations for Applied Researchers
First, we recommend that researchers present estimates of the breakdown point $\bar{r}_{X}^{\text{bp}}$ as a scalar measure of the robustness of their results. This can be displayed alongside the main result point estimates, as in Table 5.2 in our empirical application. Second, we recommend that researchers present estimates of the bounds described in Theorem 1 as a function of $\bar{r}_{X}$. Researchers may also find it useful to present estimates of the calibration parameters $\rho_{k}$ that we describe in section 3.4. We illustrate all of these recommendations in our empirical application in section 5.